# The Value of Information in Human-AI Decision-making

### v5

@@ Abstract
Multiple agents are increasingly combined to make decisions with the expectation of achieving complementary performance, where the decisions they make together outperform those made individually. However, knowing how to improve the performance of collaborating agents requires knowing what information and strategies each agent employs. With a focus on human-AI pairings, we contribute a decision-theoretic framework for characterizing the value of information. By defining complementary information, our approach identifies opportunities for agents to better exploit available information in AI-assisted decision workflows. We present a novel explanation technique (ILIV-SHAP) that adapts SHAP explanations to highlight human-complementing information. We validate the effectiveness of ACIV and ILIV-SHAP through a study of human-AI decision-making, and demonstrate the framework on examples from chest X-ray diagnosis and deepfake detection. We find that presenting ILIV-SHAP with AI predictions leads to reliably greater reductions in error over non-AI assisted decisions more than vanilla SHAP [^1] The code to calculate the main quantities in our framework and reproduce the experimental results is available at https://osf.io/p2qzy/?view_only=bf39de5d96f047f69e45ffd42689ebf9..
@@ 1 Introduction
Evidence of complementary performance between humans and AI nonetheless remains limited, with many studies showing that human-AI teams underperform the AI alone (Buçinca et al., 2020; Bussone et al., 2015; Green and Chen, 2019; Jacobs et al., 2021; Lai and Tan, 2019; Vaccaro and Waldo, 2019; Kononenko, 2001). A rigorous understanding of such results is limited by the fact that most analyses of human-AI decision-making focus on ranking the performance of human-AI teams or agents individually using measures based on post-hoc decision accuracy (Passi and Vorvoreanu, 2022). However, such approaches do not account for the best achievable performance given the information available at the time of the decision (Kleinberg et al., 2015; Guo et al., 2024; Rambachan, 2024). Moreover, existing approaches cannot provide insight on the potential for available information in the environment to improve performance. Consequently, it is difficult to design interventions that improve the team’s performance.
Imagine one could identify information complementarities that can be exploited, such as when one of the agents has access to information not contained in the other’s judgments, or has not fully integrated contextually-available information in their judgments. This would motivate interventions to improve decision-making. For example, if we can identify how much complementary information AI predictions provide over human judgments, we can use this knowledge to guide model selection or motivate further data collection to improve the model. Conversely, finding evidence that model predictions contain decision-relevant information that humans do not exploit can motivate the design of explanations communicating complementary information.
@@ 2 Related work
Many empirical studies of human-AI collaboration focus on AI-assisted human decision-making for legal, ethical, or safety reasons (Bo et al., 2021; Boskemper et al., 2022; Bondi et al., 2022; Schemmer et al., 2022). However, a recent meta-analysis by Vaccaro et al. (2024) finds that, on average, human–AI teams perform worse than the better of the two agents alone. In response, a growing body of work seeks to evaluate and enhance complementarity in human–AI systems (Bansal et al., 2021b, 2019, a; Wilder et al., 2021; Rastogi et al., 2023; Mozannar et al., 2024b). The present work differs from much of these prior works by approaching human-AI complementarity from the perspective of information value and use, including asking whether the human and AI decisions provide additional information that is not used by the other.
@@
Our work contributes methods for evaluating the decisions of human-AI teams (Kleinberg et al., 2015, 2018; Lakkaraju et al., 2017; Mullainathan and Obermeyer, 2022; Rambachan, 2024; Guo et al., 2024; Ben-Michael et al., 2024; Shreekumar, 2025). Kleinberg et al. (2015) proposed that evaluations of human-AI collaboration should be based on the information that is available at the time of the decision. According to this view, our work defines Bayesian best-attainable-performance benchmarks similar to several prior works (Guo et al., 2024; Wu et al., 2023; Agrawal et al., 2020; Fudenberg et al., 2022). Closest to our work, Guo et al. (2024) model the expected performance of a rational Bayesian agent faced with deciding between the human and AI recommendations as the theoretical upper bound on the expected performance of any human-AI team. This benchmark provides a basis for identifying exploitable information within a decision problem.
Human information in machine learning.
Some approaches focus on automating the decision pipeline by explicitly incorporating human expertise in developing machine learning models, such as by learning to defer (Mozannar et al., 2024a; Madras et al., 2018; Raghu et al., 2019; Keswani et al., 2022, 2021; Okati et al., 2021; Chen et al., 2022). Corvelo Benz and Rodriguez (2023) propose multicalibration over human and AI model confidence information to guarantee the existence of an optimal monotonic decision rule. Alur et al. (2023) propose a hypothesis testing framework to evaluate the added value of human expertise over AI forecasts. Our work shares the motivation of incorporating human expertise, but provides a general analytical framework for quantifying the information value of all available signals and agent decisions in human–AI decision tasks.
@@ 3 Framework
We cast the information available to a DM,including any available agent decisions, as a set of signals defined within an information model. Following the definition of an information model in Blackwell et al. (1951), the information model can be represented by a data-generating model with a set of signals.
  • Signals. There are $n$ “basic signals” represented as random variables $\Sigma_{1},\ldots,\Sigma_{n}$, from the signal spaces $\mathbf{\Sigma}_{1},\ldots,\mathbf{\Sigma}_{n}$. Basic signals represent information available to a decision-maker as they decide, e.g., $\mathbf{\Sigma}_{1}\in\Delta\mathbf{\Omega}$ for a probabilistic prediction of raining, $\mathbf{\Sigma}_{2}=\{\text{cloudy},\text{not cloudy}\}$, $\mathbf{\Sigma}_{3}\in\{0,\ldots,100\}$ for temperature in Celsius, etc. The decision-maker observes a signal, which is a subset of the basic signals, $V\subseteq 2^{\{\Sigma_{1},\dots,\Sigma_{n}\}}$. The combination of two signals $V_{1}$ and $V_{2}$ takes the set union $V=V_{1}\cup V_{2}$. For example, in the standard human-AI decision workflow where a human makes an independent judgment before consulting the AI, the set of all basic signals is $V=\{x,D^{\text{H}},D^{\text{AI}},e(D^{\text{AI}},x)\}$–the features of the instance, the human’s initial decisions, the AI’s predictions, and any explanation of the AI’s predictions.
@@
We adopt the same idea below to define the agent-complementary information value.
@@ 3.1 Agent-Complementary Information Value
With the above definitions, it is possible to measure the best case additional value that new signals can provide over the information already captured by an agent’s decisions. Here, agent may refer to a human, an AI system, or a human–AI team. The intuition behind our approach is that any information that is used by decision-makers should eventually reveal itself through variation in their decisions. We recover the information value in agent decisions by offering the decisions as a signal to the Bayesian rational DM.
@@
Of course, we are unlikely to observe identical realizations of the signals in continuous-valued or high-dimensional data, such as images or text. A natural relaxation is to consider the realizations that are sufficiently “similar” for the performance of the decision-maker. Algorithm 1 learns the posterior distribution of the payoff-relevant state from the signal and then uses that prediction as the probability distribution of the payoff-relevant state to choose the optimal action. This approach can be readily extended to the case where the signal is more complex, such as human decisions in the form of freeform text (e.g., radiology reports) or explanations of the AI predictions in the form of images (e.g., saliency map).
@@
1:Input: Observed realizations $\{v_{i},d^{b}_{i},\omega_{i}\}_{i=1}^{n}$, a regression algorithm $\mathcal{A}$, payoff function $S$, decision space $\mathbf{D}$, and state space $\mathbf{\Omega}$.
@@
7:  $\hat{d}^{r}_{i}\leftarrow\operatorname{argmax}_{d\in\mathbf{D}}\text{\bf E}_{\omega\sim\hat{p}_{i}}\![{S(d,\omega)}]$
@@
9:  $\hat{d}^{rb}_{i}\leftarrow\operatorname{argmax}_{d\in\mathbf{D}}\text{\bf E}_{\omega\sim\hat{p}^{b}_{i}}\![{S(d,\omega)}]$
@@ 3.2 Instance-level Agent-complementary Information Value
$\operatorname{ACIV}$ quantifies the value of the decision-relevant information in a signal $V$ across the distribution of all possible realizations defined by the data-generating model. To provide analogous instance-level quantification of information value, we define Instance-Level agent-complementary Information Value ($\operatorname{ILIV}$) to quantify the value of the decision-relevant information encoded by a single realization of the signal. This finer-grained view makes it possible to analyze how much an agent can benefit in theory from better incorporating instance-level information in their decision.
@@
where $d^{r}(v^{\prime})$ is the Bayesian optimal decision on receiving $v^{\prime}$ as defined in Equation 2. Note that we allow $v^{\prime}\neq v$, i.e. the signal $v^{\prime}$ observed by the rational DM can be different from the actual realization $v$. The expected payoff of a rational DM who is misinformed is guaranteed to be lower than is correctly informed, i.e., $\mathrm{r}^{v}(v^{\prime})\leq\mathrm{r}^{v}(v)$ for any $v^{\prime}$. This notion of flexibility allows to consider the information value of a counterfactual realization of a signal, which we use in designing the explanation of ILIV-SHAP in the next section.
@@
where $\mathrm{r}^{v}(\emptyset;D^{b})$ represents the expected rational payoff on instances of $V=v$, where the rational DM only knows the information encoded in agent decisions without knowing any realizations of $V$. $\operatorname{ILIV}$ maximizes when the signal does not misinform, i.e., $\operatorname{ILIV}^{v}(v;D^{b})\geq\operatorname{ILIV}^{v}(v^{\prime};D^{b}),\text{for }v^{\prime}\in\mathbf{V}$.
@@
1:Input: Observed realizations $\{v_{i},d^{b}_{i},\omega_{i}\}_{i=1}^{n}$, test counterfactual signal realization $v^{\prime}$, test signal realization $v$, a regression algorithm $\mathcal{A}$, payoff function $S$, decision space $\mathbf{D}$, and state space $\mathbf{\Omega}$.
@@
8:  $\hat{d}^{r}_{i}\leftarrow\operatorname{argmax}_{d\in\mathbf{D}}\text{\bf E}_{\omega\sim\hat{p}_{i}}\![{S(d,\omega)}]$
@@
10:  $\hat{d}^{rb}_{i}\leftarrow\operatorname{argmax}_{d\in\mathbf{D}}\text{\bf E}_{\omega\sim\hat{p}^{b}_{i}}\![{S(d,\omega)}]$
@@ 5 Experiment
Metric AI1 (higher ACIV) AI2 (lower ACIV)
MAPE 14.30% 14.51%
R squared 0.81 0.81
$\operatorname{ACIV}$
(in MSE)
$6.5\times 10^{8}$ $3.7\times 10^{8}$
$\operatorname{ACIV}$
(in MAPE)
4.61% 2.00%
Figure 1: Left: AI performance and ACIV. Right: Reduction in the human-AI team’s absolute percentage error (APE) with 95% confidence intervals according to posterior predictions of our regression model.
@@
Participants were assigned to one of six experimental conditions resulting from crossing the two AI models with the three explanation types. We designed AI1 and AI2 to have varying potential to complement humans: AI1 was trained with the six features and the true house price, with noise added to the six features (which is meant to be bring down the accuracy of AI1 to be comparable with AI2). On the other hand, AI2 was trained with the four human-interpretable features and the true house price, such that there would be less potential for the predictions to add complementary information over human judgments. We used the same training data for both AI models and ensured that the two AI models achieved similar performance (Figure 1 left), so that model accuracy differences would not confound our results. In the SHAP explanation, we used the SHAP values of the six features to explain the AI’s prediction. To generate the ILIV-SHAP explanations, we used independent human decisions we collected on Prolific for the dataset using the same task and features prior to running the official study. We present ILIV-SHAP explanations of the six features, which ranked and highlighted the features where the human-complementary information (ILIV) over the AI’s prediction exceeds a threshold (see Appendix H for explanation examples).
@@
We report the $\operatorname{ACIV}$ of the AI models in mean squared error (MSE) [^3] We included MSE since it is a proper scoring rule while MAPE is not. and mean absolute percentage error (MAPE). To measure the human alone and human-AI team performance, we fit a preregistered Bayesian hierarchical regression model with weakly informed priors444link to preregistration redacted for peer review to the percentage error (PE) between the human’s prediction and the true house price, $\textit{PE}=(d-\omega)/\omega$, using R’s brms package (Bürkner, 2017).

### v6

# Explaining and Improving Information Complementarities in Multi-Agent Decision-making

@@ Abstract
Multiple agents are increasingly combined to make decisions with the expectation of achieving complementary performance, where the decisions they make together outperform those made individually. However, knowing how to improve the performance of collaborating agents requires knowing what information and strategies each agent employs. With a focus on human-AI pairings, we contribute a decision-theoretic framework for characterizing the value of information. By defining complementary information, our approach identifies opportunities for agents to better exploit available information in AI-assisted decision workflows. We present a novel explanation technique (ILIV-SHAP) that adapts SHAP explanations to highlight human-complementing information. We validate the effectiveness of our framework and ILIV-SHAP through a study of human-AI decision-making, and demonstrate the framework on examples from chest X-ray diagnosis and deepfake detection. We find that presenting ILIV-SHAP with AI predictions leads to reliably greater reductions in error over non-AI assisted decisions more than vanilla SHAP [^1] The code to calculate the main quantities in our framework and reproduce the experimental results is available at https://osf.io/p2qzy/?view_only=bf39de5d96f047f69e45ffd42689ebf9..
@@ 1 Introduction
Many empirical studies, however, have found that human-AI teams underperform the AI alone (Buçinca et al., 2020; Bussone et al., 2015; Green and Chen, 2019; Jacobs et al., 2021; Lai and Tan, 2019; Vaccaro and Waldo, 2019; Kononenko, 2001). Two sources of ambiguity complicate such results. One concerns the role of measurement: performance is often scored against post-hoc decision accuracy (Passi and Vorvoreanu, 2022) rather than accounting for the best achievable performance given information available at the time of the decision (Kleinberg et al., 2015; Guo et al., 2024; Rambachan, 2024). Additionally, it often remains unclear how agents differed in their information access or use, making it difficult to design interventions to improve these aspects.
Imagine one could identify information complementarities that can be exploited, such as when one of the agents has access to information not contained in the other’s judgments, or has not fully integrated contextually-available information (e.g., instance features) in their judgments. This would motivate interventions to improve decision-making. For example, if we can identify how much complementary information AI predictions provide over human judgments, we can use this knowledge to guide model selection or motivate further data collection to improve the model. Conversely, finding evidence that model predictions contain decision-relevant information that humans do not exploit can motivate the design of explanations communicating complementary information.
@@ 2 Related work
Many empirical studies of human-AI collaboration focus on AI-assisted human decision-making for legal, ethical, or safety reasons (Bo et al., 2021; Boskemper et al., 2022; Bondi et al., 2022; Schemmer et al., 2022). However, a recent meta-analysis by Vaccaro et al. (2024) finds that, on average, human–AI teams perform worse than the better of the two agents alone. In response, a growing body of work seeks to evaluate and enhance complementarity in human–AI systems (Bansal et al., 2021b; 2019; a; Wilder et al., 2021; Rastogi et al., 2023; Mozannar et al., 2024b). The present work differs from much of these prior works by approaching human-AI complementarity from the perspective of information value and use, including asking whether the human and AI decisions provide additional information that is not used by the other.
@@
Our work contributes methods for evaluating the decisions of human-AI teams (Kleinberg et al., 2015; 2018; Lakkaraju et al., 2017; Mullainathan and Obermeyer, 2022; Rambachan, 2024; Guo et al., 2024; Ben-Michael et al., 2024; Shreekumar, 2025). Kleinberg et al. (2015) proposed that evaluations of human-AI collaboration should be based on the information that is available at the time of the decision. According to this view, our work defines Bayesian best-attainable-performance benchmarks similar to several prior works (Guo et al., 2024; Wu et al., 2023; Agrawal et al., 2020; Fudenberg et al., 2022). Closest to our work, Guo et al. (2024) model the expected performance of a rational Bayesian agent faced with deciding between the human and AI recommendations as the theoretical upper bound on the expected performance of any human-AI team. This benchmark provides a basis for identifying exploitable information within a decision problem.
Complementarity by design.
Some approaches focus on automating the decision pipeline by explicitly incorporating human expertise in developing machine learning models or human-AI collaboration pipeline, such as by learning to defer (Mozannar et al., 2024a; Madras et al., 2018; Raghu et al., 2019; Keswani et al., 2022; 2021; Okati et al., 2021; Chen et al., 2022). Corvelo Benz and Rodriguez (2023) propose multicalibration over human and AI model confidence information to guarantee the existence of an optimal monotonic decision rule. Other approaches exploit information asymmetry (i.e., cases where humans have additional contextual knowledge) and offer principled methods with provable guarantees to improve team decision performance (Straitouri et al., 2023; De Toni et al., 2024; Arnaiz-Rodriguez et al., 2025). Alur et al. (2024) propose a framework to incorporate human decisions into machine learning algorithms when the state is indistinguishable from the algorithm alone but can be discriminated by humans. Our work also concerns information asymmetry, but provides an interpretable analytical framework for quantifying the information value of all available signals and agent decisions in human–AI decision tasks, enabling the design of information-based interventions.
@@ 3 Framework
We cast the information available to a DM,including any available agent decisions, as a set of signals defined within an information model. Following the definition of an information model in Blackwell and others (1951), the information model can be represented by a data-generating model with a set of signals.
  • Signals. There are $n$ “basic signals” represented as random variables $\Sigma_{1},\ldots,\Sigma_{n}$, from the signal spaces $\mathbf{\Sigma}_{1},\ldots,\mathbf{\Sigma}_{n}$. Basic signals represent information available to a decision-maker as they decide, e.g., $\mathbf{\Sigma}_{1}=\Delta\mathbf{\Omega}$ for a probabilistic prediction of raining, $\mathbf{\Sigma}_{2}=\{\text{cloudy},\text{not cloudy}\}$, $\mathbf{\Sigma}_{3}=\{0,\ldots,100\}$ for temperature in Celsius, $\mathbf{\Sigma}_{4}=\mathbf{D}$ for human decisions $D^{H}$ on taking umbrella or not, etc. The decision-maker observes a signal, which is a subset of the basic signals, $V\subseteq 2^{\{\Sigma_{1},\dots,\Sigma_{n}\}}$. The combination of two signals $V_{1}$ and $V_{2}$ takes the set union $V=V_{1}\cup V_{2}$. In the standard human-AI decision workflow where a human makes an independent judgment before consulting the AI, all basic signals are $\{x,D^{\text{H}},D^{\text{AI}}\}$–the features of the instance, the human’s initial decisions and the AI’s predictions.
@@
The full information value in a human-AI decision task is the information value of the set of all basic signals. For example, in the weather decision example, $\operatorname{IV}(\{\Sigma_{1},\Sigma_{2},\Sigma_{3},\Sigma_{4}\})$ represents the full information value of the probabilistic prediction of rain, the cloudiness, the temperature, and the human decisions. This defines the upper bound of the information value of any signal, including the agent-complementary information value and instance-level agent-complementary information value that we will define in the following sections.
@@ 3.1 Agent-Complementary Information Value
With the above definitions, it is possible to measure the best case additional value that new signals can provide over the information already captured by an agent’s decisions. Here, agent may refer to a human, an AI system, or a human–AI team. The intuition behind our approach is that any information that is used by decision-makers should eventually reveal itself through variation in their decisions. Definition˜3.2 captures how much complementary information value is offered by a signal over the agent decisions in expectation over the data-generating process. For example, in the weather decision task, this indicates how much the payoff from the decision-maker’s choice of whether to take the umbrella or not can be further improved by incorporating the temperature in their decision rule.
@@
Of course, we are unlikely to observe identical realizations of the signals in continuous-valued or high-dimensional data, such as images or text. A natural relaxation is to consider the realizations that are sufficiently “similar” for the performance of the decision-maker. Algorithm˜1 learns the posterior distribution of the payoff-relevant state from the signal and then uses that prediction as the probability distribution of the payoff-relevant state to choose the optimal action. This approach can be readily extended to the case where the signal is more complex, such as human decisions in the form of freeform text (e.g., radiology reports) or explanations of the AI predictions in the form of images (e.g., saliency map).  [^3] Note that the algorithm $(\hat{a},\hat{a}^{b})$ should be cross-validated to avoid overfitting to the observed data and also be evaluated for calibration error since the rational DM treats it as the Bayesian posterior. We provide a sensitivity analysis in Appendix H where we compare the ACIV estimated under the rational belief estimator using linear regression (LR), gradient boosting methods (GBM), and neural networks (NN).
@@
1:Input: Observed realizations $\{v_{i},d^{b}_{i},\omega_{i}\}_{i=1}^{n}$, a predictive algorithm $\mathcal{A}$, payoff function $S$, decision space $\mathbf{D}$, and state space $\mathbf{\Omega}$.
@@
7:  $\hat{d}^{r}_{i}\leftarrow\operatorname*{arg\,max}_{d\in\mathbf{D}}\text{\bf E}_{\omega\sim\hat{p}_{i}}\![{S(d,\omega)}]$
@@
9:  $\hat{d}^{rb}_{i}\leftarrow\operatorname*{arg\,max}_{d\in\mathbf{D}}\text{\bf E}_{\omega\sim\hat{p}^{b}_{i}}\![{S(d,\omega)}]$
@@ 3.2 Instance-level Agent-complementary Information Value
$\operatorname{ACIV}$ quantifies the value of the decision-relevant information in a signal $V$ across the distribution of all possible realizations defined by the data-generating model. To provide analogous instance-level quantification of information value, we define Instance-Level agent-complementary Information Value ($\operatorname{ILIV}$) to quantify the value of the decision-relevant information encoded by a single realization of the signal. This finer-grained view makes it possible to analyze how much an agent can benefit in theory from better incorporating instance-level information in their decision. For example, in the weather decision task, this indicates how much the payoff of the decision-maker can be improved by knowing the temperature is 21C.
@@
where $d^{r}(v^{\prime})$ is the Bayesian optimal decision on receiving $v^{\prime}$ as defined in Equation˜2. Note that we allow $v^{\prime}\neq v$, i.e. the signal $v^{\prime}$ observed by the rational DM can be different from the actual realization $v$. The expected payoff of a rational DM who is misinformed is guaranteed to be lower than is correctly informed, i.e., $\mathrm{r}^{v}(v^{\prime})\leq\mathrm{r}^{v}(v)$ for any $v^{\prime}$. This notion of flexibility allows to consider the information value of a counterfactual realization of a signal. For example, in the weather decision task, this notation is able to describe how much payoff can change when the decision maker is misinformed that the temperature is 18C when the temperature is actually 21C. We use this in designing the explanation of ILIV-SHAP in the next section.
@@
where $\mathrm{r}^{v}(\emptyset;D^{b})$ represents the expected rational payoff on instances of $V=v$, where the rational DM only knows the agent decisions without knowing any realizations of $V$. $\operatorname{ILIV}$ maximizes when the signal does not misinform, i.e., $\operatorname{ILIV}^{v}(v;D^{b})\geq\operatorname{ILIV}^{v}(v^{\prime};D^{b}),\text{for }v^{\prime}\in\mathbf{V}$.
@@
1:Input: Observed realizations $\{v_{i},d^{b}_{i},\omega_{i}\}_{i=1}^{n}$, test counterfactual signal realization $v^{\prime}$, test signal realization $v$, a predictive algorithm $\mathcal{A}$, payoff function $S$, decision space $\mathbf{D}$, and state space $\mathbf{\Omega}$.
@@
8:  $\hat{d}^{r}_{i}\leftarrow\operatorname*{arg\,max}_{d\in\mathbf{D}}\text{\bf E}_{\omega\sim\hat{p}_{i}}\![{S(d,\omega)}]$
@@
10:  $\hat{d}^{rb}_{i}\leftarrow\operatorname*{arg\,max}_{d\in\mathbf{D}}\text{\bf E}_{\omega\sim\hat{p}^{b}_{i}}\![{S(d,\omega)}]$
@@ 4 Information-based Explanation
Sample-based methods for approximating SHAP values [^4] We note that ILIV-SHAP does not resolve the foundational issues of SHAP (Huang and Marques-Silva, 2024), but just as a demonstration on how the information value such as ILIV can be used in explanations. can also be applied to ILIV-SHAP, such as permutation sampling (Strumbelj and Kononenko, 2010), Kernel SHAP (Lundberg and Lee, 2017), and Partition SHAP (Lundberg et al., 2018). Because $\operatorname{ILIV}$ is expected to be non-decreasing in the number of features included in $g_{f}(\mathbf{x}^{\prime})$—more features mean more information—sample-based methods can achieve better stability across the permutations on ILIV-SHAP than on SHAP.
@@ 5 Experiment
(a) AI1 with ILIV-SHAP and SHAP explanation.
(b) AI2 with ILIV-SHAP and SHAP explanation.
Figure 1: The screenshot of the interface for ILIV-SHAP and SHAP explanations. The third column is the SHAP value and the highlights are based on the ILIV-SHAP value. Because AI2 does not have access to Feature X/Y, both their ILIV-SHAP and SHAP values are zero for AI2.
@@
[…]
@@
We report the $\operatorname{ACIV}$ of the AI models in mean squared error (MSE) [^6] We included MSE since it is a proper scoring rule while MAPE is not. and mean absolute percentage error (MAPE). To measure the human alone and human-AI team performance, we fit a preregistered Bayesian hierarchical regression model with weakly informed priors777link to preregistration redacted for peer review to the percentage error (PE) between the human’s prediction and the true house price, $\textit{PE}=(d-\omega)/\omega$, using R’s brms package (Bürkner, 2017).
@@
Table 1: The two AI models that are used in the experiment. AI1 has access to all the features, while AI2 has access to only the human-interpretable features.
Input features MAPE R squared $\operatorname{ACIV}$ (in MSE) $\operatorname{ACIV}$ (in MAPE)
AI1 All features 14.30% 0.81 6.5 $\times$ 108 4.61%
AI2 Human-interpretable features 14.51% 0.81 $3.7\times 10^{8}$ 2.00%
Figure 2: Reduction in the human-AI team’s absolute percentage error (APE) with 95% confidence intervals according to posterior predictions of our regression model.